3. Productivity and growth policies

Consider a small island country whose only industry is weaving. The following table shows information about the small economy in two different years.

Complete the table by calculating physical capital per worker as well as labor productivity.

Hint: Recall that productivity is defined as the amount of goods and services a worker can produce per hour. In this problem, measure productivity as the quantity of goods per hour of labor.

Year

Physical Capital

Labor Force

Physical Capital per Worker

Labor Hours

Output

Labor Productivity

(Looms)

(Workers)

(Looms)

(Hours)

(Garments)

(Garments per hour of labor)

2016 300 100 4,000 36,000
2017 480 120 4,200 50,400

Based on your calculations,   in physical capital per worker from 2016 to 2017 is associated with   in labor productivity from 2016 to 2017.

Suppose you're in charge of establishing economic policy for this small island country.

Which of the following policies would lead to greater productivity in the weaving industry? Check all that apply.

Sharply increasing the interest rate on student loans to people pursuing advanced degrees in weaving

Encouraging saving by allowing workers to set aside a portion of their earnings in tax-free retirement accounts

Offering free public education to every worker in the country

Imposing a tax on looms
Answer #1

(1) Physical capital per worker = Physical capital / Labor force

2016: 300/100 = 3

2017: 480/120 = 4

(2) Labor productivity = Output / Labor hours

2016: 36,000/4,000 = 9

2017: 50,400/4,200 = 12

(3) Increase in physical capital per worker is associated with Increase in labor productivity.

(4) Following policies would lead to greater productivity:

- Encouraging saving by allowing workers to set aside a portion of their earnings in tax-free retirement accounts

- Offering free public education to every worker in the country